# Critical role of phase-dependent properties in modeling photothermal sintering of LiCoO2 cathodes

### 23 Apr 2026 v1

@@ 1.1 Amorphous LCO as a proxy for as-deposited films and grain boundary regions
Figure 2: Amorphous LCO generated at different densities exhibits negligible structural differences. (a) Enthalpy and (b) density during thermal cycling under 0%, 5%, and 10% applied uniform tensile strain. Each system is heated to 4300 K for 100 ps to induce complete amorphization, rapidly quenched to 300 K over 4 ps, and then equilibrated at 300 K in the isothermal-isobaric (NpTNpT) ensemble for 1 ns to relax residual pressure. (c) Radial distribution functions $g(r)$ of the resulting amorphous structures. All three show a pronounced first peak at the characteristic metal–oxygen bond distance, dominated by Co–O correlations from local CoO6 octahedra, with a weaker Li–O contribution at a similar distance reflecting the more variable Li coordination. A broader second peak at larger $r$ arises from mixed Li–Li, Li–Co, Co–Co, and O–O correlations, and its reduced amplitude/broadening indicates only weak medium-range order. At longer distances, $g(r)$ rapidly approaches 1, confirming the absence of long-range periodicity.
@@ 3.1 Molecular dynamics simulations
To simulate melting, each of the four crystalline structures is gradually heated from 0.01 K to 3000 K over 300 ps under the isothermal-isobaric (NpTNpT) ensemble at zero pressure. Temperature and pressure control are applied using the Nose–Hoover thermostat and barostat [57], with a time step of 0.1 fs to ensure both numerical stability and physical fidelity. This setup is consistently applied in all subsequent MD simulations using Matlantis.
@@ 3.4 Calculations of LCO properties using the trained NNP
Lattice parameters, volume and density Lattice constants $a$ and $c$ for structures at 0 K are obtained by structure optimization of the 360-atom LCO using the conjugate gradient method [52] with box relaxation. The density is then calculated from the relaxed cell volume and the molar mass of LiCoO2 (97.87 g/mol). The values for 300 K are obtained by maintaining the structure at 300 K for 15 ps with atoms being relaxed under the isothermal-isobaric (NpTNpT) ensemble using the Nose-Hoover thermostat, with the temperature adjusted every 100 time steps with one integration step of 0.1 fs, and the pressure along each dimension adjusted every 1000 time steps with one integration step of 0.1 fs. The lattice constants and volume are then calculated by time-averaging over 1000 configurations for the last 5 ps. The same simulation conditions are applied to the calculations of the coefficient of thermal expansion.
@@
For each temperature, lattice parameters were first obtained from NpTNpT simulations using a timestep of 0.1 fs. The system was then thermalized in the NVTNVT ensemble for 100 ps, after which production runs were performed in the NVENVE ensemble with the heat flux recorded throughout. The ensemble average in Eq. (5) is obtained from 20 independent MD trajectories with different initial velocities. The trajectory length was adjusted for each system to ensure convergence of the HFACF, ranging from 1.6 to 2.4 ns.

### 26 Apr 2026 v2

@@ 1.1 Amorphous LCO as a proxy for as-deposited films and grain boundary regions
Figure 2: Amorphous LCO generated at different densities exhibits negligible structural differences. (a) Enthalpy and (b) density during thermal cycling under 0%, 5%, and 10% applied uniform tensile strain. Each system is heated to 4300 K for 100 ps to induce complete amorphization, rapidly quenched to 300 K over 4 ps, and then equilibrated at 300 K in the isothermal-isobaric ($NpT$) ensemble for 1 ns to relax residual pressure. (c) Radial distribution functions $g(r)$ of the resulting amorphous structures. All three show a pronounced first peak at the characteristic metal–oxygen bond distance, dominated by Co–O correlations from local CoO6 octahedra, with a weaker Li–O contribution at a similar distance reflecting the more variable Li coordination. A broader second peak at larger $r$ arises from mixed Li–Li, Li–Co, Co–Co, and O–O correlations, and its reduced amplitude/broadening indicates only weak medium-range order. At longer distances, $g(r)$ rapidly approaches 1, confirming the absence of long-range periodicity.
@@ 3.1 Molecular dynamics simulations
To simulate melting, each of the four crystalline structures is gradually heated from 0.01 K to 3000 K over 300 ps under the isothermal-isobaric ($NpT$) ensemble at zero pressure. Temperature and pressure control are applied using the Nose–Hoover thermostat and barostat [57], with a time step of 0.1 fs to ensure both numerical stability and physical fidelity. This setup is consistently applied in all subsequent MD simulations using Matlantis.
@@ 3.4 Calculations of LCO properties using the trained NNP
Lattice parameters, volume and density Lattice constants $a$ and $c$ for structures at 0 K are obtained by structure optimization of the 360-atom LCO using the conjugate gradient method [52] with box relaxation. The density is then calculated from the relaxed cell volume and the molar mass of LiCoO2 (97.87 g/mol). The values for 300 K are obtained by maintaining the structure at 300 K for 15 ps with atoms being relaxed under the isothermal-isobaric ($NpT$) ensemble using the Nose-Hoover thermostat, with the temperature adjusted every 100 time steps with one integration step of 0.1 fs, and the pressure along each dimension adjusted every 1000 time steps with one integration step of 0.1 fs. The lattice constants and volume are then calculated by time-averaging over 1000 configurations for the last 5 ps. The same simulation conditions are applied to the calculations of the coefficient of thermal expansion.
@@
For each temperature, lattice parameters were first obtained from $NpT$ simulations using a timestep of 0.1 fs. The system was then thermalized in the $NVT$ ensemble for 100 ps, after which production runs were performed in the $NVE$ ensemble with the heat flux recorded throughout. The ensemble average in Eq. (5) is obtained from 20 independent MD trajectories with different initial velocities. The trajectory length was adjusted for each system to ensure convergence of the HFACF, ranging from 1.6 to 2.4 ns.